OpenStudy (princessaurora):

*Will fan and medal* The expression BLANK is NOT equivalent to (1 − sin2(x)) tan(-x). A) (1-cos^2(x)) cot(-x) B) (cos^2(x)-1) cot(x) C) (sin^2(x)-1) tan(x) D) (cos^2(x)-1) cot(-x)

OpenStudy (johnweldon1993):

Wait, is the QUESTION supposed to be (1 - sin^2(x))(tan(-x))

OpenStudy (princessaurora):

I just copied what the question was, but maybe it can be written as that also can it?

OpenStudy (princessaurora):

Yeah what you wrote is fine.

OpenStudy (johnweldon1993):

So lets see \[\large (1 - sin^2(x))(tan(-x))\] Frst, we know tan is an odd function so \(\large tan(-x) = -tan(x)\) \[\large (1 - sin^2(x))(-tan(x))\] We know \(\large tan(x) = \frac{sin(x)}{cos(x)}\) so \[\large (1 - sin^2(x))(-\frac{sin(x)}{cos(x)})\] And finally we know \(\large 1 - sin^2(x) = cos^2(x)\) so \[\large (cos^2(x))(-\frac{sin(x)}{cos(x)})\] Leaving \[\large -cos(x)sin(x)\] Now we just need to find if any of your options simplify down to this

OpenStudy (princessaurora):

ok I see

OpenStudy (johnweldon1993):

Want me walk through it with ya or do you think you got it? :)

OpenStudy (princessaurora):

walk me through it please

OpenStudy (johnweldon1993):

Okay :) so it will all be the same exact steps A) \(\large (1-cos^2(x))(cot(-x)\) Now we know cot is also an odd function...so again \(\large cot(-x) = -cot(x)\) and we also still know that \(\large 1 - cos^2(x) = sin^2(x)\) So lets rewrite what we have so far \[\large (sin^2(x))(-cot(x))\] Now we know \(\large cot = \frac{cos(x)}{sin(x)}\) so \[\large (sin^2(x))(-\frac{cos(x)}{sin(x)})\] Leaving us with \(\large -cos(x)sin(x)\) So we know that's not the answer

OpenStudy (princessaurora):

alright I see

OpenStudy (johnweldon1993):

So go ahead and try out B and tell me what you get :)

OpenStudy (princessaurora):

I'm retarded help

OpenStudy (princessaurora):

I don't know how to type it in my calculator because it wont let me square cos or any of those

OpenStudy (johnweldon1993):

Nope you dont even need to do any of that... \[\large 1 - cos^2(x) = sin^2(x)\] and \[\large tan(x) = \frac{sin(x)}{cos(x)}\] are just trig identities, VERY important to memorize

OpenStudy (princessaurora):

ok I see

OpenStudy (johnweldon1993):

Yeah absolutely no calculator required with these :)